1) Calculate the number of moles of Cl atoms in 3.01×1024 formula units of magnesium chloride, MgCl2.

Express your answer numerically in moles.

2) In the gaseous state, chlorine exists as a diatomic molecule Cl2 (Molar mass = 70.9 g/mol). Calculate the number of moles of chlorine present in 100  g of chlorine gas.

Express the quantity in moles to three significant figures.

3) Sodium hydroxide, NaOH, is a strong base that is used in industrial synthesis and processes such as making paper.

What is the mass of 2.10 ×1022 molecules of NaOH (Molar mass = 40.0 g/mol)?

Express the mass in grams to three significant figures.

4)  Ascorbic acid, or vitamin C (C6H8O6, molar mass = 176 g/mol), is a naturally occurring organic compound with antioxidant properties. A healthy adult’s daily requirement of vitamin C is 70-90 mg. A sweet lime contains 2.86×10−4 mol of ascorbic acid.

To determine whether the ascorbic acid in a sweet lime meets the daily requirement, calculate the mass of ascorbic acid in 2.86×10−4 mol of ascorbic acid.

Express the mass in grams to three significant figures.
Answer #1

1. Moles of MgCl2 = formula units /Avogadro's number

= 3.01*10^24/ 6.02*10^23 = 5 moles

In one mole of MgCl2, there are 2 moles of Cl atoms.

In 5 moles of MgCl2, there are 5*2= 10 moles of Cl

Answer: 10 moles